OpenStudy (anonymous):

write the following Trigonometric expression in terms of sinx. csc^2x -cot^2x. not sure how to go about solving . I got this as my answer but no sure if it is right 1/sin^2x-cos^2x/sin^2x=1

OpenStudy (anonymous):

I am so lost

OpenStudy (unklerhaukus):

\[\csc^2x -\cot^2x\\=\frac1{\sin^2x}-\frac{\cos^2x}{\sin^2x}\\=\frac{1-\cos^2 x}{\sin^2x}\] now use \[\sin^2x+\cos^2x=1\qquad\implies 1-\cos^2x=\sin^2x\]

OpenStudy (anonymous):

so i was on the right track

OpenStudy (unklerhaukus):

yes,

OpenStudy (anonymous):

so it all equals to 1 ?

OpenStudy (unklerhaukus):

thats right \[\huge\color{red}\checkmark\]

OpenStudy (anonymous):

they cross cancel eachother out

OpenStudy (anonymous):

thank you for your assistance

OpenStudy (unklerhaukus):

\[=\frac{1-\cos^2 x}{\sin^2x}\\=\frac{\sin^2x}{\sin^2x}\\=\frac{\cancel{\sin^2x}}{\cancel{\sin^2x}}\\=1\] good work